OpenStudy (anonymous):

Alex designed the model of a cylindrical water tank to be installed in a park. The model has a radius of 4 cm and a height of 12 cm. The dimensions of the actual water tanks that will be installed in the park are 5 times larger than the model. The total surface area of the actual water tanks, rounded to the nearest whole number, is _________ cm2. Use 22 over 7 for pi.

OpenStudy (anonymous):

@ivettef365 @ganeshie8

ganeshie8 (ganeshie8):

start with calculating the total surface are of model cylinder

ganeshie8 (ganeshie8):

Model : r = 4 h = 12 total surface area = \(2\pi r^2 + 2\pi r h\) = ?

OpenStudy (anonymous):

one second let me calculate it

OpenStudy (anonymous):

hmm, my answer doesn't seem like it would be right :/ I got 5, 980. 55

OpenStudy (anonymous):

.44*

ganeshie8 (ganeshie8):

check again

ganeshie8 (ganeshie8):

and keep in mind, we are calculating surface area of model only. actual cylinder we will calculate later.

ganeshie8 (ganeshie8):

wolfram says 402.3 http://www.wolframalpha.com/input/?i=2*22%2F7*4%5E2%2B2*22%2F7*4*12

OpenStudy (anonymous):

hmm :/ then i'm doing something wrong

OpenStudy (anonymous):

I kept multiplying 2 x 3.14 and i'm getting 6.28

ganeshie8 (ganeshie8):

hmm ok check it later. lets trust wolfram... its super smart :)

ganeshie8 (ganeshie8):

so total surface area of model = 402.3

OpenStudy (anonymous):

okay !

ganeshie8 (ganeshie8):

since, the dimensions of actual cylinder are 5 times larger, the surface area will be 5^2 = 25 times larger.

ganeshie8 (ganeshie8):

so we multiply the model area with 25 to get the actual cylinder area : actual cylinder area = 402.3*25 = ?

OpenStudy (anonymous):

10057. 5

ganeshie8 (ganeshie8):

yeah round it to nearest whole number, it wud be : 10057

ganeshie8 (ganeshie8):

you can use wolfram as your online calculator if it helps..

OpenStudy (anonymous):

I will from now on. so the answer is 10057?

ganeshie8 (ganeshie8):

yes

OpenStudy (anonymous):

thank you ! sorry im asking for so much help

ganeshie8 (ganeshie8):

np you welcome :)